OpenStudy (anonymous):

Please Help! In a corporation, 63% of the employees are female, executives, or both. Furthermore, 54% of the employees are female, and 4% are female executives. Find the percentage of employees who are male executives.

OpenStudy (anonymous):

Let \(M\) denote male, \(F\) female, \(E\) executive and \(\overline{E}\) not executive. \[P(F\cup E)=0.63\\ P(F)=0.54\\ P(F\cap E)=0.04\] Using this info, you're asked to find \(P(M\cap E)\). Does the notation make sense?

OpenStudy (anonymous):

Alright, I understand that I have to find the intersection ,but would it be subtracting the union of the female with the intersection of the female? .63-.04?

OpenStudy (anonymous):

Not quite. There's a useful formula you'll be using: \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] Think of a Venn diagram. If you want to count everything in both \(A\) and \(B\), then you don't want to double-count any items in the sets. This is why we subtract \(P(A\cap B)\). So, the first thing to do, I think, would be to find \(P(E)\) using this formula.

OpenStudy (anonymous):

Oh wait is it subtracting .63-.54?

OpenStudy (anonymous):

I got it ! Thank You =)

OpenStudy (anonymous):

\[P(F\cup E)=P(F)+P(E)-P(F\cap E)\\ 0.63=0.54+P(E)-0.04\\ P(E)=0.13\]

OpenStudy (anonymous):

Alright I took note of this for the next problem in my h.w Thank You I truly appreciate it. I struggle a lot with probability.